At which point on the following curve does the tangent line has the largest slope ?

y=1+40x^3-3x^5


1
Expert's answer
2022-04-16T04:13:05-0400

The slope of the tangent line to the function y at the point x is equal to the derivative of this function at the given point.

"y\u2019=(1+40x^3-3x^5)\u2019=120x^2-15x^4"

Now we should find the points at which "y\u2019" has maximum.

"y\u2019\u2019=240x-60x^3=0"

"x=-2" ; "x=0" ; "x=2" .

"y’’(-3)=-720+27\cdot60>0"

"y\u2019\u2019(-1)=-240+60<0"

"y\u2019\u2019(1)=240-60>0"

"y(3)=720-27\cdot60<0"

We have two maximums "y\u2019" at the points "x=-2" and "y=2" .

"y’(-2)=y’(2)=120\cdot(-2)^2-15\cdot(-2)^4=240"

Answer: "(-2,240)" and "(2,240)" .